OpenStudy (anonymous):

Simplify this expression. 4 power8 ÷ 4 power4

OpenStudy (solomonzelman):

you can use ^ (shift + 6) to write "power" . \(\Huge\color{blue}{ \sf \frac{4^{8} }{4^{4}}=4^{8-4}=? }\)

OpenStudy (anonymous):

wait is that a divided sign crap in that case you just subtract the powers so it is 4^4

OpenStudy (anonymous):

\[\large \frac{ 4^8 }{ 4^4 }=\frac{4*4*4*4*4*4*4*4}{4*4*4*4}=4*4*4*4=4^{8-4}\]